A friend or yours is considering two cell phone service providers. Provider A charges $100 per month for the service regardless of the number of phone calls made. Provider B does not have a fixed service ree but instead charges $1 per minute ror calls. Your friends monthly demand for minutes or calling is given by the equation p-120-30P, where P is the price of a minute. with Provider A, the cost of an extra minute is 5 with Provider B, the cost of an extra minute is 5 Given your friends demand for minutes and the cost of an extra minute with each provider, if your friend used Provider A, he would talk for minutes, and if he used Provider B, he would talk for This means your friend would pay S for service with Provider A and for service with Provider B Use the folmowing graph to draw your friends demand curve for minutes. Then use the green triangle to help you answer the questions that follow. Note: You will not be graded on any changes you make to the graph 5.00 Demand 3.50 Triangle 2.00 20 40 0 100 120 140 160 180 200 Quantity af MinutesYour friend would obtain S in consumer surplus with Provider B. Given this information, which provider would you recommend that your friend choose? O Provider A O Provider B

A friend of yours is considering two cell phone service providers. Provider A charges $100 per month for the service regardless of the number of phone calls made. Provider B does not have a fixed service ree but instead charges $1 per minute for calls. Your friend's monthly demand for minutes or calling is given by the equation p-120-30P, where P is the price of a minute. 

With Provider A, the cost of an extra minute is 5 _______ .with Provider B, the cost of an extra minute is _______ .

Given your friend's demand for minutes and the cost of an extra minute with each provider, if your friend used Provider A, he would talk for _______ minutes, and if he used Provider B, he would talk for _______ minutes.

This means your friend would pay _______  for service with Provider A and _______ for service with Provider B.

Use the following graph to draw your friend's demand curve for minutes. Then use the green triangle to help you answer the questions that follow.

 Note: You will not be graded on any changes you make to the graph


Your friend would obtain _______  in consumer surplus with provider and _______ in consumer surplus with Provider B. 

Given this information, which provider would you recommend that your friend choose? 
Answer #1

Demand curve : Q = 120-30*p

So inverse demand curve : 30*p = 120-Q

P= 4 - Q/30

So maximum willingness to pay for minutes of calling is $4

So with A, cost of extra minute is 0 $,for B , it is $1.

since provider A is charging a fixed fees irrespective of the number of calls, so it's MC of extra unit of calling is zero.hence it charges price equals to zero.

So with A, friend will buy Q=120,

& with B, he will talk for 90 minutes,

( As, p=1, Q = 120-30 = 90 )

Thus payment to A is $ 100 & to B, payment is $90

Now graph :

n A 4 AX(4-1):90 9。120

With A, Friend will obtain consumer surplus equals to .5*120*4 = $240

From B, C.S. = .5*90*3 = $135

Thus MCQ:

Option A) since more Consumer surplus with provider A)